OpenStudy (anonymous):

Simplify the logarithm log2(32)

OpenStudy (anonymous):

\[y = \log_{b}x \rightarrow x = b^{y}\]

OpenStudy (anonymous):

log(5) I think... not sure

OpenStudy (anonymous):

When doing logarithms, the first thing you wanna do is factor.

OpenStudy (anonymous):

\[ 32 = 2^5 \]

OpenStudy (anonymous):

\[ \log_2(2^5) = 5\log_2(2) \]

OpenStudy (anonymous):

Once you factor you can use various properties: \[ \log(a\times b) = \log(a) + \log(b) \\ \log(a^b) = b\times \log (a)\\ \log_a(a) = 1 \]